# A posteriori closure of turbulence models: are symmetries preserved ?

### v1

@@ 4 Results
In this section, our DL-closed reduced system of equations, hereon referred to as LES-NN, is evaluated by comparing its predictions against the fully resolved ground truth across multiple statistical measures. We analyze its performance in reproducing key turbulence characteristics, including structure functions, temporal evolution of energy, and scale-dependent correlations. Particular attention is given to the ability of the model to capture the tails of joint probability density functions between resolved and unresolved variables; as well as scale invariant statistics following the third hypothesis of K62 theory, and whether this symmetry is preserved by the closure.
@@
$S_{n}^{(p)}=\langle|u_{n}|^{p}\rangle\,.$ (4)
These are shown in Figure 3 up to the cutoff shell, for $p=1,\ldots,6$ for both the fully resolved system (ground truth) and the LES-NN model. There is a great agreement between the model and the ground truth for all order structure functions showed, even close to the cutoff scale, which is the subgrid scales effect is more noticeable.
@@
In Figure 4, we show the time evolution of the logarithm of the normalised velocity field (velocity field scales as $u_{n}\!\sim\!\smash{k_{n}^{-1/3}}$) for both the ground truth (GT) and the LES-NN model, along with the normalized error between the two. Both the GT and LES-NN models start from the same initial condition (IC). It is evident that the LES-NN model rapidly decorrelates from the GT. This behavior is expected, as we cannot synchronize past the Lyapunov time. For this particular IC, the discrepancy is particularly clear because we have chosen an IC that is within a burst of energy that has already propagated to the small scales, where the dynamics are much faster. As a result, the effective Lyapunov time aligns with these faster time scales.
Figure 4: Time evolution of the normalised energy density $\log_{2}(|u_{n}|^{2}k_{n}^{2/3})$ for (a) the ground truth and (b) the LES-NN model. (c) Absolute normalised error $\left||u_{n}|^{2}-|\tilde{u}_{n}|^{2}\right|k_{n}^{2/3}$ between the two fields. The x-axis represents the wavenumber index $n$, while the y-axis represents the normalized time $t/\tau_{0}$. The dashed line represents the cutoff scale.
Figure 5 shows the joint PDF between the sum of the resolved energy of the shells used as input for the neural network, ${\tilde{u}_{N_{c}},\tilde{u}_{N_{c}-1},\tilde{u}_{N_{c}-2}}$, and the sum of the unresolved energy, i.e., the sum of the energies of the shells that the NN outputs, ${\tilde{u}^{\theta}_{N_{c}+1},\tilde{u}_{N_{c}+2}^{\theta}}$, for both the ground truth and the LES-NN model. While the overall shape of the PDF is correctly predicted by the model, it appears to primarily capture the mean behavior rather than accurately reproducing the full distribution.
There are a few potential reasons for this discrepancy. One possible cause is the architecture of the neural network itself (multi-layer perceptron), which may lack the necessary complexity to fully capture the higher-order correlations in the data. Spectral bias, a tendency of neural networks to favor learning smooth, low-frequency components over more complex high-frequency variations, could further contribute to the underestimation of the tails in the predicted distribution. Additionally, it can also be an outcome of the choice of the loss function as an MSE between velocity fields of the resolved shells, thus not including the unresolved ones explicitly.
@@
where $\delta_{i}v_{j}(x,\ell)=v_{j}(x+\ell\mathbf{e}_{i})-v_{j}(x)$ represents the increment of the $j$-th velocity component across a spatial separation $\ell$ along the unit vector $\mathbf{e}_{i}$. Kolmogorov posited that at sufficiently high Reynolds numbers, the statistical distributions of these multipliers would exhibit universality—dependent solely on the scale ratio $\ell/\ell^{\prime}$ and independent of the absolute scales $\ell$ or $\ell^{\prime}$. Furthermore, he hypothesized statistical independence between multipliers associated with scales separated by large intervals, a conjecture later termed his third hypothesis [14].
@@
$\Delta_{n}(t)=\theta_{n}(t)-\theta_{n-1}(t)-\theta_{n-2}(t)\,,$ (7)
where $\theta_{n}=\arg u_{n}$.
In Figure 6, we show the PDF of the amplitude and phase multipliers for the ground truth and our model, LES-NN, for shells $n\!=\!6,\ldots,N_{c}$. For the fully resolved model, we see the expected collapse. However, for the LES-NN model, the learned closure breaks this symmetry, and the multipliers are no longer scale invariant near the cutoff. This is a limitation of the model.
Figure 6: PDF of the (a) amplitude multipliers, $w_{n}$, and (b) phase multipliers, $\Delta_{n}$, for the ground truth (GT) and LES-NN model, for the shells $7$ to $N_{c}$.
We extend this analysis in Figure 7 by examining the joint PDF of the amplitude of the product of the last two resolved multipliers and the product of the two unresolved multipliers (which include the neural network’s predictions). The LES-NN model accurately captures the resolved multipliers but fails significantly in the unresolved ones, leading to incorrect predictions not only in the tails but also in the mean. As in all highly nonlinear, high-dimensional optimization problems, the neural network training converged to a local minimum—sufficient to reproduce high-order structure functions correctly but learning a flawed way to match the subgrid-scale statistics, resulting in good overall results but with clear failures in certain observables. This suggests room for improvement, possibly in the choice of loss function, input-output representation, or overall problem formulation. We leave this as an open question for future investigation.
@@
Following Eyink et al. [16], we can further define:
$\sigma_{n}(t)=\ln w_{n}(t)\,,$ (8)
$U_{n}(t)=\exp(i\Delta_{n}(t))\,,$ (9)
where $\sigma_{n}(t)$ is a “local slope” of the multiplier $w_{n}(t)$ and $U_{n}(t)$ is a 2D rotator spin that lives on the unit circle in the complex plane. We are interested in studying the correlation of these variables across different shells. For this purpose, we define the following correlation function [16]:
$C_{XY}(n,m)=\left\langle X_{n}^{*}Y_{m}\right\rangle-\left\langle X_{n}^{*}\right\rangle\left\langle Y_{m}\right\rangle\,.$ (10)
@@
The results for $X\!=\!Y\!=\!\sigma$ and $X\!=\!Y\!=\!U$ shown in Figure 8 for the GT and LES-NN model. Shell $n\!=\!11$ is kept as the reference to compute correlations. We see the expected fast decay with increasing $|m-n|$; however, the correlations near the cutoff are wrongly estimated for the LES-NN model. This helps shine a light on the lack of collapse of the multipliers near the cutoff.
Figure 8: Spin-spin correlation functions with shell $n\!=\!11$ as reference. The dashed line represents the cutoff scale.
5 Discussion
In this work, we have delved deeper into the performance of a closure trained using a solver-in-the-loop approach for closure of the shell models of turbulence [1]. We focused on correlations between resolved and unresolved variables, as well as universal single-time statistics of the inertial range of fully developed turbulence, the so-called Kolmogorov multipliers (ratio between adjacent shells). We showed that, even though our closure is able to reproduce high-order moments, such has structure functions of $p=6$ order with great accuracy, it still underpredicts the cross-correlations between the energy of the resolved and unresolved shells. While the overall shape of the joint PDF is well captured, the learned closure tends to focus on the mean behavior. This could be attributed to several factors. One potential reason is the architecture of the neural network in use (i.e., a simple multi-layer perceptron), which might lack the complexity needed to fully capture higher-order correlations. Additionally, the spectral bias inherent in neural networks could cause them to prioritize smoother, low-frequency components, leading to an underestimation of the tails in the predicted distribution. The choice of the loss function, which focuses solely on the MSE between velocity fields of the resolved shells, may also contribute, as it does not explicitly account for the unresolved shells.
Regarding the performance of the learned closure in terms of the multipliers, the LES-NN model breaks their scale invariance near the cutoff: the PDFs of the multipliers near the cutoff no longer collapse. The perfect LES model would preserve this symmetry in the inertial range. Our model was not trained using multipliers, so it is not completely surprising. The interesting question is whether training the model focusing on multiplier statistics in the loss function will yield a better-performing model than one trained on simple velocity differences. This was explored by Domingues Lemos et al. in [7] and more recently in [8] by working in the hidden symmetry framework [9]. Compared to these, our closure performs better on observables such as structure functions and flatness, but not in terms of multipliers statistics. Training with a solver-in-the-loop approach and within the multipliers framework might result in a better-learned closure. Another powerful benefit of using the multipliers closure is that, in principle, the learned closure should be cutoff independent (given that the cutoff is not affected by the forcing or the dissipation scales). This was numerically demonstrated by Domingues Lemos in [8].
We report these results in the hope of advancing our understanding of turbulence closure problems, particularly in relation to the optimal setup, which has strong implications for both data-driven based and phenomenology based closures. Additionally, the study on multiplier statistics, and the evidence that our model breaks this symmetry are potential signs that the current framework is not ideal, and one that explicitly contains or enforces this scale invariance might be more attractive. This has potential implication in subgrid scale modeling in 3D Navier-Stokes turbulence, as a similar rescaling procedure can be done, leading to scale invariance [17]. Naturally, additional challenges are introduced, but the overall message might still remain true. Future studies on the topic are, therefore, needed.

### v2

@@ 4 Results
In this section, our DL-closed reduced system of equations, hereon referred to as LES-NN, is evaluated by comparing its predictions against the fully resolved ground truth across multiple statistical measures. We analyze its performance in reproducing key turbulence characteristics, including structure functions, temporal evolution of energy, and scale-dependent correlations. Particular attention is given to the ability of the model to capture the tails of joint probability density functions between resolved and unresolved variables; as well as scale invariant statistics following the third hypothesis of K62 theory [23], and whether this symmetry is preserved by the closure. All LES-NN results are obtained starting from initial conditions not used during training, and the trajectories are evolved for time horizons substantially longer than those present in the training data (details in Appendix A). Thus, the form of generalization assessed here is in-distribution: robustness to unseen initial conditions and to long-time evolution, rather than out-of-distribution extrapolation across different Reynolds numbers or cutoff scales.
@@
$S_{n}^{(p)}=\langle|u_{n}|^{p}\rangle,$ (6)
where $\langle\cdot\rangle$ denotes averaging in time and across different initial conditions. These are shown in Figure 3 up to the cutoff shell, for $p=1,\ldots,6$ for both the fully resolved system (ground truth) and the LES-NN model. There is a great agreement between the model and the ground truth for all orders of structure functions showed, even close to the cutoff scale, which is where the effect of the subgrid scales is more noticeable.
@@
In Figure 4, we show the time evolution of the logarithm of the normalised energy field (velocity field scales as $u_{n}\!\sim\!\smash{k_{n}^{-1/3}}$) for both the ground truth (GT) and the LES-NN model, along with the normalized error between the two. Both the GT and LES-NN models start from the same initial condition (IC). It is evident that the LES-NN model rapidly decorrelates from the GT. This behavior is expected, as we cannot synchronize past the Lyapunov time. For this particular IC, the discrepancy is particularly clear because we have chosen an IC that is within a burst of energy that has already propagated to the small scales, where the dynamics are much faster. As a result, the effective Lyapunov time for the LES can be roughly approximated as the time scale of the cutoff shell.
Figure 4: Time evolution of the normalised energy density $\log_{2}(|u_{n}|^{2}k_{n}^{2/3})$ for (a) the ground truth and (b) the LES-NN model. (c) Absolute normalised error $\left||u_{n}|^{2}-|\tilde{u}_{n}|^{2}\right|k_{n}^{2/3}$ between the two fields. The x-axis represents the wavenumber index $n$, while the y-axis represents time normalised by the integral time scale $t/\tau_{0}$. The dashed line represents the cutoff scale.
In order to assess the cross-correlations between resolved and unresolved scalesFigure 5 shows the joint PDF between the summed energy of the resolved shells used as input to the neural network, $\tilde{u}_{N_{c}},\tilde{u}_{N_{c}-1},\tilde{u}_{N_{c}-2}$}, and the sum of the unresolved energy, i.e., the sum of the energies of the shells that the NN outputs, $\tilde{u}^{\theta}_{N_{c}+1},\tilde{u}_{N_{c}+2}^{\theta}$}, for both the ground truth and the LES-NN model. The overall shape of the distribution is well captured, but deviations appear in the joint tails, indicating that the learned closure underestimates the strength of the cross-scale correlations, which are connected with the correlation of the shell phases. We discuss possible causes of this discrepancy at the end of this section.
@@
where $\delta_{i}v_{j}(x,\ell)=v_{j}(x+\ell\mathbf{e}_{i})-v_{j}(x)$ represents the increment of the $j$-th velocity component across a spatial separation $\ell$ along the unit vector $\mathbf{e}_{i}$. Kolmogorov posited that at sufficiently high Reynolds numbers, the statistical distributions of these multipliers would exhibit universality–dependent solely on the scale ratio $\ell/\ell^{\prime}$ and independent of the absolute scales $\ell$ or $\ell^{\prime}$. Furthermore, he hypothesized statistical independence between multipliers associated with scales separated by large intervals, a conjecture later termed his third hypothesis [7].
@@
$\Delta_{n}(t)=\varphi_{n}(t)-\varphi_{n-1}(t)-\varphi_{n-2}(t)\,$, (9)
where $\varphi_{n}=\arg u_{n}$.
In Figure 6, we show the PDF of the amplitude and phase multipliers for the ground truth and our model, LES-NN, for shells $n\!=\!7,\ldots,N_{c}$. For the fully resolved model, we observe the expected collapse. However, for the LES-NN model, the learned closure breaks this symmetry, and the multipliers are no longer scale invariant near the cutoff. This is a limitation of the model.
Figure 6: PDF of the (a) amplitude multipliers, $w_{n}$, and (b) phase multipliers, $\Delta_{n}$, for the ground truth and LES-NN model, for the shells 77 to $N_{c}$.
We extend this analysis in Figure 7 by examining the joint PDF of the amplitude of the product of the last two resolved multipliers and the product of the two unresolved multipliers (which include the neural network’s predictions). The LES-NN model accurately captures the resolved multipliers but fails significantly in the unresolved ones, leading to incorrect predictions not only in the tails but also in the mean.
@@
To further probe correlations, following Eyink et al. [14], we introduce variables analogous to the spins of the two-dimensional XY model [3, 24],
$\sigma_{n}(t)=\ln w_{n}(t),\qquad U_{n}(t)=\exp\!\big(i\,\Delta_{n}(t)\big),$ (10)
where $\sigma_{n}$ quantifies local amplitude fluctuations and $U_{n}$ encodes relative phase differences between neighboring shells. The set $\{U_{n}\}$ behaves as interacting rotators on the unit circle, whose alignment measures the degree of phase coherence across scales. In this analogy, the energy cascade corresponds to the propagation of order in a spin chain: strong alignment implies coherent phase dynamics and efficient energy transfer. To quantify this behavior, we compute correlation functions,
$C_{XY}(n,m)=\langle X_{n}^{*}Y_{m}\rangle-\langle X_{n}^{*}\rangle\langle Y_{m}\rangle,$ (11)
@@
for $X\!=\!Y\!=\!\sigma$ and X=Y=UX\!=\!Y\!=\!U, which measure how amplitude and phase coherence decay with scale separation. The reference shell, n=11n\!=\!11, is chosen within the inertial range, close to but below the cutoff. The results are shown in Figure 8 for the GT and LES-NN model. We see the expected fast decay with increasing |mn||m-n|; however, the correlations near the cutoff are wrongly estimated for the LES-NN model. This helps shine a light on the lack of collapse of the multipliers near the cutoff.
Figure 8: Correlation functions $C_{XY}(n,m)$ for amplitude fluctuations ($X\!=\!Y\!=\!\sigma$) and phase rotations (X=Y=UX\!=\!Y\!=\!U), using shell n=11n\!=\!11 as the reference within the inertial range. The curves show how correlations decay as the shell separation |mn||m-n| increases for both the ground truth and the LES–NN model. The dashed vertical line marks the cutoff shell $N_{c}$. While both models display the expected rapid decay at large separations, the LES–NN correlations deviate near the cutoff, indicating a loss of phase and amplitude coherence in this region.
The results presented above highlight both the strengths and limitations of the learned closure. The LES-NN model reproduces the resolved-scale statistics with high accuracy, including the overall scaling of the structure functions up to high orders, although small deviations appear at the highest orders. However, these residual differences become more evident when we examine observables that explicitly couple resolved and unresolved variables, such as the joint PDFs and the multipliers near the cutoff. In these quantities, the model underestimates the tails of the distributions and fails to recover the full joint variability between scales. This analysis thus acts as a magnifying glass, revealing the origin of the small but systematic departures observed in the higher-order statistics of the resolved variables.
Several factors could in principle contribute to these discrepancies. One possibility is the neural-network architecture itself: a fully connected MLP may not possess sufficient expressive power to capture all higher-order correlations, and spectral bias could suppress the small-scale variability associated with strong intermittency. However, in our tests, changing the network architecture and increasing its depth or width (above what we considered for the results presented here) did not lead to significant improvements. Similarly, changing the loss function to include the subgrid shells explicitly, rather than only the resolved variables, did not alter the results. These observations suggest that the origin of the discrepancies does not lie primarily in architectural choices or in the omission of subgrid shells from the loss function.
Instead, our findings point to a structural limitation arising from the formulation of the closure problem itself, particularly the Markovian nature of the model. The Mori–Zwanzig (MZ) formalism [37, 45] provides a natural perspective on this issue. When a high-dimensional dynamical system is projected onto a reduced set of resolved variables (as it is done in LES), the resulting exact evolution equation contains two additional contributions beyond the Markovian term: a memory kernel, encoding the delayed interactions between resolved and unresolved variables, and an orthogonal dynamics term that accounts for unresolved-unresolved interactions that cannot be reproduced with the memory term and are usually modeled as noise. A purely Markovian closure omits both effects, and thus misses key ingredients that should in principle help in capturing the correlations between resolved and unresolved variables. This structural lack of memory offers a possible explanation for the breakdown of scale-invariant multiplier statistics near the cutoff and the incomplete reconstruction of resolved-unresolved correlations observed in our results.
Possible ways forward include extending the framework to incorporate non-Markovian effects in the spirit of the MZ formulation, for instance by learning effective memory-kernel or orthogonal-dynamics terms directly from data, as recently explored in data-driven MZ approaches [31, 11]. In parallel, one may also improve the closure by explicitly incorporating the known scale invariance of multipliers into the learning process, either as a penalty in the loss function or as hard constraint on the model architecture, so that the collapse of the multiplier PDFs is enforced rather than left as an emergent property. Another alternative is to reformulate the problem in terms of multipliers themselves by rescaling the governing equations, as done by Domingues Lemos et al. [12, 28] within the hidden-symmetry framework [34]. Combining these ideas—for example, by training a solver-in-the-loop model directly in the multiplier variables or by enforcing multiplier symmetries within the present formulation—may lead to closures that simultaneously reproduce high-order resolved statistics and respect the scale-invariant structure of the multipliers. These considerations are an attempt at clarifying the origin of the observed discrepancies and point to concrete directions for improving the closure framework.
5 Conclusion
In this work, we have examined in detail a data-driven closure for a shell model of turbulence trained with a solver-in-the-loop (a posteriori) strategy [16]. Building on our previous study, which focused mainly on high-order statistics of the resolved variables, we have gone beyond single-scale observables and analysed quantities that explicitly couple resolved and unresolved degrees of freedom. In particular, we investigated joint probability density functions between resolved and subgrid shells, Kolmogorov multipliers and their scale invariance, and correlation functions inspired by the XY-model analogy [14, 3, 24].
Our results show that our data-driven closure reproduces resolved-scale statistics with high accuracy, including structure functions up to order p=6p\!=\!6, and remains stable under long-time integration. However, when we probe different observables that evaluate the correlations between resolved and unresolved shells, systematic discrepancies emerge. The joint PDFs between resolved and unresolved energies reveal an underestimation of the tails, indicating that the model does not fully capture the variability across the cutoff. Similarly, the amplitude and phase multipliers lose their expected scale invariance near the cutoff, and the spin–spin correlations involving shells close to the cutoff are slightly misestimated. These diagnostics act as a magnifying glass: they expose limitations that are not apparent from the resolved-scale statistics alone.
We have argued that these shortcomings are unlikely to be resolved merely by increasing network capacity or by including subgrid shells directly in the loss function, and instead point to a possible structural limitation of a Markovian closure. Within the Mori-Zwanzig formalism [37, 45], an exact coarse-grained description necessarily involves a memory kernel and an orthogonal dynamics term. A memoryless closure, such as the one considered here, cannot fully reproduce these effects, which offers a plausible explanation for the breakdown of multiplier scale invariance near the cutoff and the incomplete reconstruction of resolved-unresolved correlations.
These findings suggest several directions for future work. On the one hand, incorporating non-Markovian effects in a principled way, for example via data-driven MZ operators [31, 11], may allow closures that better capture multi-scale correlations across the cutoff. On the other hand, explicitly embedding key symmetries of turbulence into the learning problem—such as the scale invariance of multipliers, or working directly in the multiplier/hidden-symmetry formulation [12, 28, 34]—could lead to models that are both more universal and less sensitive to the choice of cutoff. Extending and testing these ideas in three-dimensional Navier–Stokes turbulence, where similar rescaling procedures can be defined [33], remains an important challenge. Overall, our analysis suggests that combining solver-in-the-loop training with a physically informed structure is a promising route towards robust and better-performing data-driven subgrid scale closures.
@@ Acknowledgments
Appendix A Training details
The numerical simulations and solver-in-the-loop training procedure were performed using the parameters listed in Table 1. The corresponding training loop is summarized in Algorithm 1. The algorithm represents a single training iteration of the solver-in-the-loop scheme. At each time step, the neural network predicts the unresolved shells, $\tilde{u}^{>,\theta}_{t}$, from the resolved state $\tilde{u}^{<}_{t}$. The resolved dynamics are then advanced in time, and the closure terms $\mathcal{C}_{N_{c}-1}$ and $\mathcal{C}_{N_{c}}$, which correspond to the contributions excluded from the direct integration because they depend on the predicted unresolved shells, are subsequently added through explicit integration. This procedure is repeated for msteps iterations, producing trajectories used to evaluate the multi-step loss that measures the deviation from the reference solution. Gradients of the loss with respect to the neural-network parameters are then obtained via backpropagation, and the parameters are updated using gradient-based optimization. All code developed for this work is openly available at github.com/andremfreitas/sm_loop.
Table 1: Values of the parameters used in the numerical experiments.
Parameter Value Description
$\nu$ $1\times 10^{-12}$ Viscosity
$\mathrm{Re}$ $\approx 10^{12}$ Reynolds number
$\epsilon$ $0.5$ Forcing amplitude
NN $40$ Number of shells
$N_{\eta}$ $30$ Kolmogorov scale
$N_{\mathrm{c}}$ $14$ Subgrid cutoff scale
$\tau_{0}$ $7.553\times 10^{-1}$ Eddy turnover time at the integral scale
$\tau_{\eta}$ $1.8367\times 10^{-6}$ Eddy turnover time at the dissipative scale
$\Delta t$ $1\times 10^{-8}$ Timestep of ground truth simulation
$\Delta\tilde{t}$ $1\times 10^{-5}$ Timestep of LES–NN model
$N_{\mathrm{data}}$ $256$ Number of initial conditions in the dataset
$N_{\mathrm{batch}}$ $1024$ Batch size used for training
$T_{\mathrm{train}}$ $1.65\,\tau_{0}$ Integration time of training dataset
$T_{\mathrm{test}}$ $3.31\,\tau_{0}$ Integration time of test dataset
msteps $100$ Time steps in-the-loop per training iteration
Algorithm 1 Training loop (single training iteration).
1:Initialize gradient tape $\triangleright$ Record operations for differentiation
[…]
3:for $t=0$ to msteps1msteps-1 do
[…]
6:  $\mathcal{C}_{N_{c}-1}\leftarrow\Delta\tilde{t}\,i(ak_{N_{c}}\tilde{u}^{\theta}_{N_{c}+1}\tilde{u}^{*}_{N_{c}})$ $\triangleright$ Compute closure term for shell $N_{c}-1$
7:  $\mathcal{C}_{N_{c}}\leftarrow\Delta\tilde{t}\,i(ak_{N_{c}+1}\tilde{u}^{\theta}_{N_{c}+2}\tilde{u}^{*,\theta}_{N_{c}+1}+bk_{N_{c}}\tilde{u}^{\theta}_{N_{c}+1}\tilde{u}^{*}_{N_{c}-1})$ $\triangleright$ Compute closure term for shell $N_{c}$
8:  $\mathcal{C}\leftarrow\text{concatenate}(\mathcal{C}_{N_{c}-1},\mathcal{C}_{N_{c}})$ $\triangleright$ Aggregate closure corrections
9:  $\tilde{u}^{<}_{t+1}\leftarrow\tilde{u}^{<}_{t+1}+\mathcal{C}$ $\triangleright$ Update resolved shells with closure contribution
10:end for
11:Compute loss $\mathcal{L}$ $\triangleright$ Compare predicted and reference trajectories
12:Compute gradients $\nabla_{\theta}\mathcal{L}$ $\triangleright$ Differentiate loss w.r.t. network parameters
13:Update parameters $\theta$ $\triangleright$ Apply optimizer step